If the store currently charges a price of $50, then increases that price to $60, what happens to total revenue from shoe sales (calculate P * Q before and after the price change)? Repeat the exercise for initial prices being decreased to $40 and $20, respectively. 


1
Expert's answer
2022-01-10T09:55:43-0500


If the store charges $50, then the revenue "= 50 \times 200 = \$10000"

If the store charges $60, then the revenue "= 60 \times 100 = \$6000"

The total revenue decreases in this case

Now, when the price decreases to $40, then the total revenue is "= 40 \times 300 = \$12000"

Further reducing the price to $20, the total revenue is "= 20 \times 500 = \$10000"